OpenStudy (anonymous):

-13(-45a+27)=13(-8a-10)

OpenStudy (aroub):

Distribute!

OpenStudy (aroub):

-13(-45a+27)=13(-8a-10) 585a-351=-104a-130 585a+104a=351-130 689a=221 a=17/53

OpenStudy (anonymous):

It's 1 :)

OpenStudy (aroub):

o.O

OpenStudy (kymber):

Distribute Add 351 to both sides Add 104A to both sides Divide by 689 A = .32075471698... Hope this helps! :]